A long horizontal wire contains a current such that 9.0✕1018 electrons per second flow through any given point from left to right. Determine the (a)current and (b)magnitude of the field of the long wire at a point 10 mm parallel to it in the north.


1
Expert's answer
2021-05-16T19:37:02-0400

a.) "6.25 \times 10^{18}" electrons per second gives current "= 1A"


Hence, "9 \times 10^{18}" electrons gives current "= \dfrac{9 \times 10^{18}}{6.25 \times 10^{18}} = 1.44A"


b.) Magnetic field can be given as,


"B = \dfrac{\mu_0i}{2 \pi d}"



"B = \dfrac{1.25 \times 10^{-6} \times 1.44}{2 \times 3.14 \times 10 \times 10^{-3}}"


"B = 0.28 \times 10^{-2}T"